# See What I Mean? CUE: A Cognitive Model of Understanding Explanations

### v1

@@ 1 Introduction
Machine Learning (ML) systems are increasingly used to support high-stakes decisions. Consequently, the need for human-understandable explanations has become a central concern in the field of Explainable AI (XAI) [15, 26, 3, 32]. However, most current evaluation methods focus on technical properties such as faithfulness [48, 2], robustness [38, 6], or semantical properties such as polysemanticity or clarity [17, 16]. Although these methods consider crucial aspects of XAI evaluation, they are not addressing visual properties, impacting how well users can perceive, process, and make sense of the explanations presented [30, 14]. This challenge becomes especially critical as explanations should be accessible to a diverse spectrum of users, including those with visual impairments and other cognitive disabilities [42, 54]. Given this disconnect, we focus in this work on understanding how we understand visual explanations.
@@
In the context of our work, we define visual explanations as any information presented to users with the intention of explaining a machine learning model’s inference, provided that it can be perceived via the human visual system. This includes images, figures, and text, excluding non-visual modalities (e.g. audio) [5].
@@
To better understand how users engage with visual explanations, we turn to the concept of readability, originally developed in linguistics to describe “what makes some texts easier to read than others” [18]. Foundational work [27, 28] showed how visual factors like character size and contrast affect reading speed, while others [12] framed readability as factors affecting readers’ success with printed material.
@@
A crucial distinction has emerged in linguistics between legibility, the visual clarity of text (e.g., font size, layout, contrast), and readability, the cognitive ease of understanding what is read [53, 22]. A document with large font size may be visually legible but difficult to read due to dense or overly technical language. These insights have extended beyond text to domains such as code comprehension, interface design, and data visualization [57, 52, 43].
@@
As the research of data visualization has evolved, a field concerned with optimizing visualization of complex datasets for a broad audience [8], the study of readability and legibility has expanded beyond linguistics to the realm of graphical interpretation. In the field of data visualizations, legibility refers to the clarity of individual visual elements (such as bars or lines), while readability involves understanding the overall narrative conveyed by these elements [19, 10]. Together, these concepts play a critical role in minimizing cognitive load and maximizing the effectiveness of a visualization by facilitating easier interpretation.
@@
Explainable AI visualizations, such as heatmaps [56] and concept visualizations [1], can be seen as a specialized form of data visualization. Similar to traditional data visualizations, they rely on visual encodings to communicate complex information, but they introduce unique challenges related to interpreting model behavior. While there has been a growing body of research directed to designing XAI, it has primarily focused on the general interface and interaction design [55, 29, 11, 37]. We suggest that findings from data visualization research, particularly those not only concerning with optimizing layout and interaction design but also with specific visual attributes such as color intensity, contrast, or brightness, offer valuable insights for improving the understanding of explanations.
@@ 2 Related Work
Layer-wise Relevance Propagation (LRP) [4], Local Interpretable Model-agnostic Explanations (LIME) [46], and Gradient-weighted Class Activation Mapping (Grad-CAM) [50] are common methods in XAI for compiling attributions. LRP and Grad-CAM are model-specific, using internal structures (e.g., gradients), while LIME is model-agnostic, relying on local perturbations [51].
Heatmaps are an essential visualization technique in XAI, using color gradients to depict continuous, non-binary feature attribution, activation, attention, or saliency across activation tensors [20]. Formally defined as $n$-dimensional spatial projections of model-derived attribution scores ($n\geq 2$), heatmaps are commonly applied in the image domain to highlight which input features contribute to model predictions [40]. In detection tasks, they localize specific elements within complex images, while in classification, they help users assess whether predictions rely on consistent and task-relevant regions. Techniques for generating heatmaps include sensitivity-based methods, deconvolution, and LRP [48, 49, 47].
Recent frameworks, such as PREVis [10], have made strides toward formalizing understanding within the domain of data visualization. PREVis assesses understandability, layout clarity, and data value/pattern readability. These frameworks aim to assess and improve how well users can comprehend complex visualizations, offering valuable insights to improve the design of visual elements and their interpretation.
A recent contribution by [51] introduces an abilities-based framework for understanding in XAI, distinguishing six capacities: recognizing, assessing, predicting, intervening, explaining, and designing. This approach emphasizes tailoring explanations to context-specific stakeholder needs. In the given context, that could be applicants assessing fairness versus developers designing systems. Unlike traditional monolithic approaches to XAI, the framework integrates insights from philosophy, psychology, and computer science, linking user abilities to goals such as fairness and trust.
Whilst [51] emphasize the abilities users need to interact with AI explanations, our model focuses on the process of understanding. We show how sensory perception (legibility) evolves into causal reasoning (interpretability) through structured stages. Speith provides a foundation for the actions users can take based on explanations, while our model outlines how users cognitively reach that point. Together, these frameworks offer complementary insights for XAI design, aligning ethical and functional goals with a cognitive pathway.
@@ 3 CUE: A formalized Cognitive Model of Understanding Explanations
Figure 1: Adapted from the work of Cabouat et al. [10], we propose CUE, a cognitive model for understanding explanations and integrate it with our definition of three properties of XAI understandability: legibility, readability and interpretability.
According to [9], reading a visualization requires transforming raw visual features into meaningful internal representations, a process we extend to the understanding of explanations. Building on prior cognitive models of visualization comprehension [10], we propose a structured framework for modeling the Cognitive Understanding of Explanations — CUE.
@@ 3.1 The CUE Model in Detail
At the center of CUE is the XAI algorithm, which generates an output, the explanation, designed to make the ML model’s inference understandable to the user. Once created, the explanation is presented through an explanation display, which can take various forms such as verbal, visual, or diagrammatic representations of the explanation [31]. The choice of display plays a critical role in how effectively individuals can perceive and engage with the explanation [39, 13, 44, 24].
@@
  • Good Example: A counterfactual explanation: "Loan denied because income is \$5k below threshold. Increase to \$50k for approval."
  • Poor Example: A vague statement: "Model output reflects input patterns," offering no causal insight.
@@ 3.2 Influencing Factors: The Display, the Goals and the Individual’s Knowledge
The task or context provides individuals with goals that direct their attention during sensory processing and shape their perception of relevant visual features. When reviewing an explanation, individuals operate within a specific task context, aiming to achieve certain goals [39]. These objectives can be divided into two categories: Explanation goals and task goals. Explanation goals refer to the individual’s purpose related to the explanation itself, such as evaluating the accuracy of a model’s prediction or enhancing its performance [35]. Task goals, on the other hand, define the desired outcomes for task completion, such as minimizing time or effort (see Figure 2).
The individual’s knowledge significantly influences both the comprehension and interpretation of the explanation. Visualization knowledge refers to the individual’s familiarity with display conventions, such as understanding axes, colormaps, and other visualization techniques, often referred to as “visualization literacy” [7]. Domain knowledge refers to the individual’s expertise in the subject matter of the visualization, including the underlying algorithm, the model being explained, the data, and any additional context needed to interpret the explanation [21].
@@ 4.1 Experiment Design
We selected image classification with convolutional neural networks (CNN) and heatmap explanations as a proxy task modality as it represents one of the most foundational and widely studied applications in explainable AI. Proxy tasks are controlled experimental setups that simplify real-world explanation scenarios and allow one to isolate variables (e.g., colormap design), control task complexity, and ensure replicability [30]. In addition, this modality allowed us to control task complexity, standardize stimuli, and ensure replicability, which is essential for an initial investigation.
@@
Figure 5: Example of LRP explanation displayed with colormaps BWR, Coolwarm and Cividis.
To address the research questions, we selected three colormaps: BWR, Coolwarm (common but not accessibility-optimized) [34] from Matplotlib [23], and Cividis, which was specifically designed for accessibility by users with color vision deficiency [41]. Participants with and without visual impairments evaluated heatmap explanations across these conditions, completing tasks to measure:
@@ 5.2 RQ2: Accessibility-Focused Colormaps and Usability Gaps
Contrary to expectations, Cividis worsened confidence ($r=0.462$) and efficiency ($r=0.524$) disparities compared to BWR ($r=0.254$), despite its perceptual optimizations (Figure 5). Task performance remained unchanged ($p>0.05$), suggesting that the presence of “accessible” design elements does not necessarily lead to more accessible XAI systems.
@@ 7 Conclusion
Also, the findings underscore the need for explanation systems that go beyond visual optimization. Adaptive, multi-modal interfaces, allowing for interactive and tailored engagement, may better support diverse user needs [36, 33, 45, 25], especially for individuals with limited visual capacity.
@@
The Ethics Commission of the Fraunhofer Heinrich Hertz Institute provided guidelines for the study procedure and determined that no protocol approval is required. Informed consent has been obtained from all participants.

### v2

@@ 1 Introduction
Machine Learning (ML) systems are increasingly used to support high-stakes decisions. Consequently, the need for human-understandable explanations has become a central concern in the field of Explainable AI (XAI) Doshi-Velez and Kim (2017); Lage et al. (2019); Arrieta et al. (2020); Longo et al. (2024). However, most current evaluation methods focus on technical properties such as faithfulness Samek et al. (2017); Agarwal and Nguyen (2020), robustness Montavon et al. (2018); Bhatt et al. (2020), or semantical properties such as polysemanticity or clarity Dreyer et al. (2024, 2025). Although these methods consider crucial aspects of XAI evaluation, they are not addressing visual properties, impacting how well users can perceive, process, and make sense of the explanations presented Liao and Varshney (2021); Donoso-Guzmán et al. (2023). This challenge becomes especially critical as explanations should be accessible to a diverse spectrum of users, including those with visual impairments and other cognitive disabilities Nwokoye et al. (2024); Tielman et al. (2024). Given this disconnect, we focus in this work on understanding how we understand visual explanations.
@@
In the context of our work, we define visual explanations as any information presented to users with the intention of explaining a machine learning model’s inference, provided that it can be perceived via the human visual system. This includes images, figures, and text, excluding non-visual modalities (e.g. audio) Becker et al. (2024).
@@
To better understand how users engage with visual explanations, we turn to the concept of readability, originally developed in linguistics to describe “what makes some texts easier to read than others” DuBay (2004). Foundational work Legge et al. (1985, 1987) showed how visual factors like character size and contrast affect reading speed, while others Dale and Chall (1949) framed readability as factors affecting readers’ success with printed material.
@@
A crucial distinction has emerged in linguistics between legibility, the visual clarity of text (e.g., font size, layout, contrast), and readability, the cognitive ease of understanding what is read Tekfi (1987); Hossain and Banerjee (2024). A document with large font size may be visually legible but difficult to read due to dense or overly technical language. These insights have extended beyond text to domains such as code comprehension, interface design, and data visualization Zuffi et al. (2007); Strizver (2013); Oliveira et al. (2020).
@@
As the research of data visualization has evolved, a field concerned with optimizing visualization of complex datasets for a broad audience Brodlie et al. (2012), the study of readability and legibility has expanded beyond linguistics to the realm of graphical interpretation. In the field of data visualizations, legibility refers to the clarity of individual visual elements (such as bars or lines), while readability involves understanding the overall narrative conveyed by these elements Franconeri et al. (2021); Cabouat et al. (2024). Together, these concepts play a critical role in minimizing cognitive load and maximizing the effectiveness of a visualization by facilitating easier interpretation.
@@
Explainable AI visualizations, such as heatmaps Zhou et al. (2016) and concept visualizations Achtibat et al. (2023), can be seen as a specialized form of data visualization. Similar to traditional data visualizations, they rely on visual encodings to communicate complex information, but they introduce unique challenges related to interpreting model behavior. While there has been a growing body of research directed to designing XAI, it has primarily focused on the general interface and interaction design Wolf (2019); Liao et al. (2020); Chromik and Butz (2021); Mohseni et al. (2021). We suggest that findings from data visualization research, particularly those not only concerning with optimizing layout and interaction design but also with specific visual attributes such as color intensity, contrast, or brightness, offer valuable insights for improving the understanding of explanations.
@@ 2 Related Work
Layer-wise Relevance Propagation (LRP) Bach et al. (2015), Local Interpretable Model-agnostic Explanations (LIME) Ribeiro et al. (2016), and Gradient-weighted Class Activation Mapping (Grad-CAM) Selvaraju et al. (2017) are common methods in XAI for compiling attributions. LRP and Grad-CAM are model-specific, using internal structures (e.g., gradients), while LIME is model-agnostic, relying on local perturbations Speith et al. (2024).
Heatmaps are an essential visualization technique in XAI, using color gradients to depict continuous, non-binary feature attribution, activation, attention, or saliency across activation tensors Gu (2022). Formally defined as $n$-dimensional spatial projections of model-derived attribution scores ($n\geq 2$), heatmaps are commonly applied in the image domain to highlight which input features contribute to model predictions Nauta et al. (2023). In detection tasks, they localize specific elements within complex images, while in classification, they help users assess whether predictions rely on consistent and task-relevant regions. Techniques for generating heatmaps include sensitivity-based methods, deconvolution, and LRP Samek et al. (2017); Selvaraju et al. (2020); Richard et al. (2021).
While several studies have evaluated attribution techniques using heatmap visualizations, they have been criticized for insufficient consideration of cognitive processes Lopes et al. (2022) or for limited generalizability due to a narrow focus on specific explanation techniques or evaluation methods Kim et al. (2024). In contrast, Kares et al. (2025) proposed a broader evaluation framework that integrates subjective, objective, and computational measures.
Recent frameworks, such as PREVis Cabouat et al. (2024), have made strides toward formalizing understanding within the domain of data visualization. PREVis assesses understandability, layout clarity, and data value/pattern readability. These frameworks aim to assess and improve how well users can comprehend complex visualizations, offering valuable insights to improve the design of visual elements and their interpretation.
A recent contribution by Speith et al. (2024) introduces an abilities-based framework for understanding in XAI, distinguishing six capacities: recognizing, assessing, predicting, intervening, explaining, and designing. This approach emphasizes tailoring explanations to context-specific stakeholder needs. In the given context, that could be applicants assessing fairness versus developers designing systems. Unlike traditional monolithic approaches to XAI, the framework integrates insights from philosophy, psychology, and computer science, linking user abilities to goals such as fairness and trust.
Whilst Speith et al. (2024) emphasize the abilities users need to interact with AI explanations, our model focuses on the process of understanding. We show how sensory perception (legibility) evolves into causal reasoning (interpretability) through structured stages. Speith provides a foundation for the actions users can take based on explanations, while our model outlines how users cognitively reach that point. Together, these frameworks offer complementary insights for XAI design, aligning ethical and functional goals with a cognitive pathway.
@@ 3 CUE: A formalized Cognitive Model of Understanding Explanations
Figure 1: Adapted from the work of Cabouat et al. Cabouat et al. (2024), we propose CUE, a cognitive model for understanding explanations and integrate it with our definition of three properties of XAI understandability: legibility, readability and interpretability.
According to Cabouat et al. (2023), reading a visualization requires transforming raw visual features into meaningful internal representations, a process we extend to the understanding of explanations. Building on prior cognitive models of visualization comprehension Cabouat et al. (2024), we propose a structured framework for modeling the Cognitive Understanding of Explanations — CUE.
@@ 3.1 The CUE Model in Detail
At the center of CUE is the XAI algorithm, which generates an output, the explanation, designed to make the ML model’s inference understandable to the user. Once created, the explanation is presented through an explanation display, which can take various forms such as verbal, visual, or diagrammatic representations of the explanation Lim et al. (2023). The choice of display plays a critical role in how effectively individuals can perceive and engage with the explanation Narayanan et al. (2018); Dhanorkar et al. (2021); Poursabzi-Sangdeh et al. (2021); Kaur et al. (2022).
@@
  • Good Example: A counterfactual explanation following the structure of Dai et al. (2022): ”If your income had been \$50k, your loan would have been approved.”
  • Poor Example: A vague statement: ”Model output reflects input patterns,” offering no causal insight.
@@ 3.2 Influencing Factors: The Display, the Goals and the Individual’s Knowledge
The task or context provides individuals with goals that direct their attention during sensory processing and shape their perception of relevant visual features. When reviewing an explanation, individuals operate within a specific task context, aiming to achieve certain goals Narayanan et al. (2018). These objectives can be divided into two categories: Explanation goals and task goals. Explanation goals refer to the individual’s purpose related to the explanation itself, such as evaluating the accuracy of a model’s prediction or enhancing its performance Meske et al. (2022). Task goals, on the other hand, define the desired outcomes for task completion, such as minimizing time or effort (see Figure 2).
The individual’s knowledge significantly influences both the comprehension and interpretation of the explanation. Visualization knowledge refers to the individual’s familiarity with display conventions, such as understanding axes, colormaps, and other visualization techniques, often referred to as “visualization literacy” Boy et al. (2014). Domain knowledge refers to the individual’s expertise in the subject matter of the visualization, including the underlying algorithm, the model being explained, the data, and any additional context needed to interpret the explanation Hegarty (2011).
@@ 4.1 Experiment Design
We selected image classification with convolutional neural networks (CNN) and heatmap explanations as a proxy task modality as it represents one of the most foundational and widely studied applications in explainable AI. Proxy tasks are controlled experimental setups that simplify real-world explanation scenarios and allow one to isolate variables (e.g., colormap design), control task complexity, and ensure replicability Liao and Varshney (2021). In addition, this modality allowed us to control task complexity, standardize stimuli, and ensure replicability, which is essential for an initial investigation.
@@
To address the research questions, we selected three colormaps: BWR, Coolwarm (common but not accessibility-optimized) Matplotlib Developers (2025) from Matplotlib Hunter (2007), and Cividis, which was specifically designed for accessibility by users with color vision deficiency Nuñez et al. (2018). Participants with and without visual impairments evaluated heatmap explanations across these conditions, completing tasks to measure:
@@ 5.2 RQ2: Accessibility-Focused Colormaps and Usability Gaps
Contrary to expectations, Cividis worsened confidence ($r=0.462$) and efficiency ($r=0.524$) disparities compared to BWR ($r=0.254$), despite its perceptual optimizations. Task performance remained unchanged ($p>0.05$), suggesting that the presence of “accessible” design elements does not necessarily lead to more accessible XAI systems.
@@ 7 Conclusion
Also, the findings underscore the need for explanation systems that go beyond visual optimization. Adaptive, multi-modal interfaces, allowing for interactive and tailored engagement, may better support diverse user needs Miller (2019); Madumal et al. (2019); Raees et al. (2024); Labarta et al. (2024), especially for individuals with limited visual capacity.
@@
The Ethics Commission of the Fraunhofer Heinrich-Hertz-Institute provided guidelines for the study procedure and determined that no protocol approval is required. Informed consent has been obtained from all participants.